# Principled RLHF from Heterogeneous Feedback via Personalization and Preference Aggregation

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) has been an effective technique for aligning AI systems with human values, with remarkable successes in fine-tuning large-language models recently. Most existing RLHF paradigms make the underlying assumption that human preferences are relatively homogeneous, and can be encoded by a single reward model. In this paper, we focus on addressing the issues due to the inherent heterogeneity in human preferences, as well as their potential strategic behavior in providing feedback. Specifically, we propose two frameworks to address heterogeneous human feedback in principled ways: personalization-based one and preference-aggregation-based one. For the former, we propose two approaches based on representation learning and clustering, respectively, for learning multiple reward models that trades-off the bias (due to preference heterogeneity) and variance (due to the use of fewer data for learning each model by personalization). We then establish sample complexity guarantees for both approaches. For the latter, we aim to adhere to the single-model framework, as already deployed in the current RLHF paradigm, by carefully aggregating diverse and truthful preferences from humans. We propose two approaches based on reward and preference aggregation, respectively: the former utilizes both utilitarianism and Leximin approaches to aggregate individual reward models, with sample complexity guarantees; the latter directly aggregates the human feedback in the form of probabilistic opinions. Under the probabilistic-opinion-feedback model, we also develop an approach to handle strategic human labelers who may bias and manipulate the aggregated preferences with untruthful feedback. Based on the ideas in mechanism design, our approach ensures truthful preference reporting, with the induced aggregation rule maximizing social welfare functions.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has been widely used in fine-tuning Large Language Models like ChatGPT and Claude 3, showing their significance and usefulness in aligning these models with human values/utilities that can be nuanced and complicated (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). The underlying assumption of RLHF is that human utility functions are closely linked to human preferences, making it feasible to learn about human utility from preference data. Therefore, most of the RLHF use a single utility function (also known as a reward function) to model the preferences of human labelers (Ziegler et al., 2019). Contrary to the assumption of “homogeneity” in reward valuation, humans assign “heterogeneous” reward values to the same question-and-answer pairs, especially for sensitive and open-ended questions, depending on their background. This diversity suggests that the probability of selecting a particular answer cannot be universally proportional to a presumed common reward value (Figure 1). Therefore, the assumption that having a homogeneous preference for humans, which is arguably an overly simplified model could potentially lead to reward models that prioritize the preferences of the majority group in the data while neglecting the preferences of minorities. This is concerning because it can result in biases and inequalities in the fine-tuned LLM systems, making it less fair and representative of diverse perspectives that are inherent in society.
@@
In addressing the challenge of diverse human preferences, the notion of personalized reward models and LLMs emerges: Intuitively, one could consider creating tailored LLMs for each human user who provides their own personalized data. However, this method might be inadequate due to the limited data available for each individual, potentially leading to inaccurate reward estimations and reduced precision in LLMs due to high variance. Conversely, employing a single reward and associated language model for every human user by naively pooling all users’ data often relies on an implicit assumption that humans have homogeneous preferences, which is barely the case in practice (Pollak and Wales, 1992; Boxall and Adamowicz, 2002).
On the other hand, even when acknowledging the heterogeneity of human preferences and aggregating them carefully to learn a single reward model, there is a challenge that has not received enough attention in the RLHF literature: humans are by nature rational (of certain degrees) and strategic, with their own objectives to optimize. They might manipulate their own feedback by not providing truthful one, to distort the aggregated results to be closer to their own preferences, and thus manipulate the output of LLMs fine-tuned over the aggregated preference data. For example, in online rating systems, users may provide extreme feedback to disproportionately influence the overall ratings toward their viewpoint.
Given the challenges above, we raise and attempt to address two questions: (1) how should we mitigate the variance of the personalized reward model and LLM model arising from the paucity of each human user’s data, and (2) if we use a single LLM model, as already deployed in the current RLHF pipeline, what constitutes an effective method for aggregating diverse and truthful preferences for fine-tuning LLMs?
Contributions.
We provide a systematic study to address the challenges above due to the heterogeneity of human preferences in RLHF, with provable efficiency guarantees. We develop two frameworks: one based on personalization, and one based on human preference aggregation. The personalization-based framework seeks to address the question (1) above, and includes two approaches as detailed below:
  • Representation-learning-based Personalization: We propose a representation learning-based approach for learning the heterogeneous individual reward function (see Figure 1). By leveraging comparison data from a diverse set of humans, we can enhance the accuracy of the representation learning (Theorem 3.4). This improvement results in better sample complexity for estimating the personalized reward model and, consequently, more effective policy learning compared to approaches using non-diverse but heterogeneous data sets (Theorem 3.1). To improve the sample complexity, we develop a new proof technique to bound the reward estimation errors without accurate representations, leading to the use of summations of non-zero expectation random variables, which may be of independent interest (proof of Lemma 5).
  • Clustering-based Personalization: We then consider personalization with human user clustering, with sample complexity guarantees. We introduce an approach for learning clustered reward functions, and personalizing users’ reward model within each cluster by the learned model for that cluster.
The aggregation-based framework seeks to address question (2), which aggregates heterogeneous human preferences in a principled way via social choice theory, and includes two aggregation approaches as detailed below:
  • Reward Aggregation with Comparison Data: We first estimate the parameters for each individual’s reward model using the individual’s preference comparison data. Then, we aggregate reward models using a family of reward aggregation rules, including those based on the utilitarianism and Leximin approaches. We then provide sample complexities of the policy induced from the single aggregated reward model.
  • Preference Aggregation with Probabilistic Opinion Data: We then provide aggregation methods for RLHF with a new feedback form: probabilistic opinion (Dietrich and List, 2016), where human labelers provide feedback as a probability vector instead of binary comparison data, as in the existing RLHF pipeline. Without assuming the relationship between human reward and preference, we directly aggregate the diverse probabilistic preferences into a consensus one. Under this feedback model, we also develop a mechanism to handle strategic human labelers, who may benefit by reporting untruthful preferences, while biasing the aggregated preference and thus the fine-tuned LLM. Our mechanism guarantees that the labelers will report truthful preferences, while the induced aggregation rules maximize various social welfare functions.
1.1 Related Work
@@
Empirical evidence has demonstrated the efficacy of incorporating human preferences into reinforcement learning (RL) for enhancing robotics (Abramson et al., 2022; Hwang et al., 2023) and for refining large-scale language models (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). These human inputs take various forms, such as rankings (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022), demonstrations (Finn et al., 2016), and scalar ratings (Warnell et al., 2018). A few approaches have been explored empirically to personalize RLHF. For example, assigning fine-grained rewards to small text segments to enhance the training process (Wu et al., 2024), or training each human labeler’s reward model with Multi-Objective Reinforcement Learning perspective (Jang et al., 2023; Hwang et al., 2023) have been proposed. Moreover, (Li et al., 2024) suggested the training of each human labeler’s reward model directly using personalized feedback with human embedding obtained by the human model, and also an approach for the clustering with finding cluster embedding.
On the theory front, the studies of RLHF have received increasing research interest. The most related prior works are (Zhu et al., 2023; Zhan et al., 2023; Wang et al., 2024), where (Zhu et al., 2023) investigated the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952) within the context of a linear reward framework; while (Zhan et al., 2023) generalized the results to encompass more general classes of reward functions. Both works concern the setting with offline preference data. (Xiong et al., 2023) provided a theoretical analysis for KL-regularized RLHF. In the online setting, (Wang et al., 2024) established a correlation between online preference learning and online RL through a preference-to-reward interface. Yet, to the best of our knowledge, there is no prior work that has analyzed RLHF with heterogeneous feedback with theoretical guarantees (except the recent independent works discussed in detail below).
@@
Early work of (Baxter, 2000) established a generalization bound that hinges on the concept of a task generative model within the representation learning framework. More recently, (Tripuraneni et al., 2021; Du et al., 2020) demonstrated that, in the setup with linear representations and squared loss functions, task diversity can significantly enhance the efficiency of learning representations. Moreover, (Tripuraneni et al., 2020) provided a representation learning with general representation and general loss functions. Representation learning has been extended to the reinforcement learning setting as well. For low-rank Markov Decision Processes, where both the reward function and the probability kernel are represented through the inner products of state and action representations with certain parameters, (Agarwal et al., 2020; Ren et al., 2022; Uehara et al., 2021) explored the theoretical foundations for learning these representations. Also, (Ishfaq et al., 2024; Bose et al., 2024) analyzed the sample complexity of multi-task offline RL.
@@
Preference aggregation is the process by which multiple humans’ preference orderings of various social alternatives are combined into a single, collective preference or choice (List, 2013). Arrow’s Impossibility Theorem demonstrates that no aggregation rule for preference orderings can simultaneously meet specific criteria essential for ensuring a fair and rational aggregation of each human user’s preferences into a collective decision (Arrow, 1951). Therefore, people considered replacing preference orderings with assigning real numbers to social alternatives (Sen, 2018; Moulin, 2004), which is sometimes called a reward (welfare) function §§§In our paper, we regard the reward function as a welfare function in social choice theory. for each human user. (Skiadas, 2016; Moulin, 2004) provided reward (welfare) aggregation rules which satisfy several desirable properties. Furthermore, an alternative method to circumvent Arrow’s impossibility theorem involved aggregating preferences via probabilistic opinion (Stone, 1961; Lehrer and Wagner, 2012). In this approach, opinions are represented as probability assignments to specific events or propositions of interest.
@@
More recently, (Zhong et al., 2024) provided a theoretical analysis of reward aggregation in RLHF, focusing primarily on linear representations. Our work, in comparison, considers general representation functions and general relationships between reward function and preference. Unlike (Zhong et al., 2024), where they focused on reward aggregation, we focus on personalization for every human labeler and also employ clustering techniques for personalization. (Zhong et al., 2024) and our paper also both investigated the case that reward and preference are not related. Our paper suggested a probabilistic opinion pooling with a mechanism design to effectively elicit truthful human preferences, presuming human labelers may be strategic. In contrast, (Zhong et al., 2024) analyzed an algorithm for a von Neumann winner policy, where a von Neumann winner policy is a policy that has at least a 50% chance of being preferred compared to any other policy. Moreover, (Zhong et al., 2024) also explored the Pareto efficiency of the resulting policy.
Notation.
The matrix $\mathbf{O}$ denotes an all-zero matrix, while $I$ stands for an identity matrix, of proper dimensions. We use $A\succ\mathbf{O}$ to denote that matrix $A$ is a positive definite matrix. The function $\sigma$ represents the Sigmoid function, defined by $\sigma(x)=1/(1+\exp(-x))$. The notation $[K]$ denotes the set $\{1,2,\dots,K\}$. $\Delta(\mathcal{A})$ refers to a probability vector in $\mathbb{R}^{|\mathcal{A}|}$. The term $\sigma_{k}^{2}(A)$ denotes the $k$-th largest singular value of matrix $A$. A function $f(x)$ is categorized based on the complexity notation as follows: $f(x)=O(g(x))$ if there exists $C>0$ and $x_{0}$ such that $f(x)\leq Cg(x)$ holds for all $x\geq x_{0}$, $f(x)=\Omega(g(x))$ if there exists $C>0$ and $x_{0}$ such that $f(x)\geq C\cdot g(x)$ for all $x\geq x_{0}$, $f(x)=o(g(x))$ if $\lim_{x\to\infty}\frac{f(x)}{g(x)}=0$, and $f(x)=\widetilde{O}(g(x))$ if $f(n)=O(g(x)\cdot\log^{k}(g(x)))$ for some finite $k$. The vector $e_{1}$ is defined as the standard basis vector of proper dimension with the first component being $1$. For a finite-dimensional vector $x$, the norm $\|x\|_{1}$ refers to its $\ell_{1}$-norm, while $\|x\|$ refers to the $\ell_{2}$-norm, unless otherwise specified. We also define $\|x\|_{\Sigma}=\sqrt{x^{\intercal}\Sigma x}$ for a positive definite matrix $\Sigma$. For a matrix $M$, the norm $\|M\|_{F}$ denotes the Frobenius norm of $M$. The multinomial distribution is denoted by $\text{Multinomial}(p_{1},\ldots,p_{n})$, where $p_{1},\ldots,p_{n}$ are the probabilities of outcomes for each of the $n$ categories, respectively, with $\sum_{i=1}^{n}p_{i}=1$ and $p_{i}\geq 0$ for all $i\in[n]$. Kullback-Leibler (KL) divergence between two probability distributions $P,Q\in\Delta(X)$ is defined as $\sum_{x\in\text{supp}(X)}P(x)\log\left(\frac{P(x)}{Q(x)}\right)$.
2 Preliminaries
Most existing RLHF processes (for language model fine-tuning) consist of two main stages: (1) learning a model of human rewards (oftentimes from preference data), and (2) fine-tuning with the reference policy through Reinforcement Learning algorithms, e.g., Proximal Policy Optimization (PPO) (Schulman et al., 2017). It may also be possible to avoid the explicit learning of reward functions while fine-tuning the policy directly from preference data (Rafailov et al., 2024). We will introduce both approaches in detail below, and will seek to develop frameworks that will cover both cases.
Markov Decision Processes.
We define the state $s$ as an element of the set of possible prompts or questions, denoted by $\mathcal{S}$, and the set of actions $a$, contained in $\mathcal{A}$, as the potential answers or responses to these questions. Consider an RLHF setting with $N$ human labelers (or users), each of whom has their own reward function. This setting can be characterized by a Markov Decision Process (MDP) with $N$ reward functions, represented by the tuple $M=(\mathcal{S},\mathcal{A},H,(P_{h})_{h\in[H]},\bm{r}=(r_{i})_{i\in[N]})$, where $H$ denotes the length of the horizon, $P_{h}:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is the state transition probability at step $h\in[H]$, ${\mathcal{T}}:=(\mathcal{S}\times\mathcal{A})^{H}$ denotes the set of all possible trajectories, and $r_{i}:{\mathcal{T}}\rightarrow\mathbb{R}$ is the reward function for individual $i$ and trajectory $\tau\in{\mathcal{T}}$, representing the utility of human user $i$ from a sequence of responses to a given prompt. We assume $-R_{\max}\leq r_{i}(\tau)\leq R_{\max}$ for every $\tau\in{\mathcal{T}}$ and $i\in[N]$, for some $R_{\max}>0$. This reward model also covers the case that $r_{i}(\tau)=\sum_{h\in[H]}r_{h,i}(s_{h},a_{h})$, where $r_{h,i}:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ denotes the state-action reward function for each step $h$ and individual $i$, and $\tau=(s_{1},a_{1},s_{2},a_{2},\dots,s_{H},a_{H})$. The MDP concludes at an absorbing termination state with zero reward after $H$ steps.
A policy $\pi_{h}:(\mathcal{S}\times\mathcal{A})^{h-1}\times{\mathcal{S}}\to\Delta(\mathcal{A})$ is defined as a function mapping trajectories to distributions over actions for each step $h\in[H]$ within the horizon $H$. We define the history-dependent policy class as $\Pi$. The collection of these policies across all steps is denoted by $\pi$. The expected cumulative reward of a policy $\pi$ is given by $J(\pi;r_{i}):=\mathbb{E}_{\tau,\pi}[r_{i}(\tau)]$ where the expectation in the formula is taken over the distribution of the trajectories under the policy $\pi$. Trajectory occupancy measures, denoted by $d_{\pi}:{\mathcal{T}}\to[0,1]$, are defined as $d_{\pi}(\tau):=\mathbb{P}_{\pi}(\tau)$, which denotes the probability of generating trajectory $\tau$ following policy $\pi$.
Relationship between Preference and Reward Function.
For the MDP with $M=(\mathcal{S},\mathcal{A},H,(P_{h})_{h\in[H]},\bm{r}=(r_{i})_{i\in[N]})$, if we compare two trajectories $\tau_{0}$ and $\tau_{1}$, we define some random variable $o$ such that $o=0$ if $\tau_{0}\succ\tau_{1}$, and $o=1$ if $\tau_{0}\prec\tau_{1}$. Here, $\tau_{0}\succ\tau_{1}$ indicates that $\tau_{0}$ is preferred than $\tau_{1}$. We assume that $P_{r_{i}}(o=0\mid\tau_{0},\tau_{1})=\Phi(r_{i}(\tau_{0})-r_{i}(\tau_{1}))$ for all $i\in[N]$, where $\Phi:\mathbb{R}\to[0,1]$ is a monotonically increasing link function, which satisfy $\Phi(x)+\Phi(-x)=1$ and $\log\Phi(x)$ is a strongly convex function. For example, $\Phi(x)=\sigma(x)$ indicates the BTL model (Definition 2.1 below), a frequently used model for the relationship between preference and reward. Also, we define $P_{\bm{r}}(\cdot\mid\tau_{0},\tau_{1}):=(P_{r_{1}}(\cdot\mid\tau_{0},\tau_{1})^{\intercal},\dots,P_{r_{N}}(\cdot\mid\tau_{0},\tau_{1})^{\intercal})^{\intercal}$. We call $P_{\bm{r}}$ and $P_{r_{i}}$ a preference probability vector induced by the reward vector $\bm{r}$ and the reward $r_{i}$.
Definition 2.1.
Direct Preference Optimization (DPO) (Rafailov et al., 2024).
[…]
@@ 2 Preliminaries
3 Personalized RLHF via Representation Learning

### v2

# RLHF from Heterogeneous Feedback via Personalization and Preference Aggregation

@@ Abstract
Reinforcement learning from human feedback (RLHF) has been an effective technique for aligning AI systems with human values, with remarkable successes in fine-tuning large-language models recently. Most existing RLHF paradigms make the underlying assumption that human preferences are relatively homogeneous, and can be encoded by a single reward model. In this paper, we focus on addressing the issues due to the inherent heterogeneity in human preferences, as well as their potential strategic behavior in providing feedback. Specifically, we propose two frameworks to address heterogeneous human feedback in principled ways: personalization-based one and preference-aggregation-based one. For the former, we propose two approaches based on representation learning and clustering, respectively, for learning multiple reward models that trade-off the bias (due to preference heterogeneity) and variance (due to the use of fewer data for learning each model by personalization). We then establish sample complexity guarantees for both approaches. For the latter, we aim to adhere to the single-model framework, as already deployed in the current RLHF paradigm, by carefully aggregating diverse and truthful preferences from humans. We propose two approaches based on reward and preference aggregation, respectively: the former utilizes social choice theory to aggregate individual reward models, with sample complexity guarantees; the latter directly aggregates the human feedback in the form of probabilistic opinions. Under the probabilistic-opinion-feedback model, we also develop an approach to handle strategic human labelers who may bias and manipulate the aggregated preferences with untruthful feedback. Based on the ideas in mechanism design, our approach ensures truthful preference reporting, with the induced aggregation rule maximizing social welfare functions.
@@ 1 Introduction
As AI models are becoming more powerful, there is greater emphasis on aligning their performance and priorities with the preferences of human users. In this context, reinforcement learning from human feedback (RLHF) has emerged as a promising approach, because it combines pre-trained large language models with direct human feedback (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). RLHF utilizes human feedback in the form of preferences over multiple responses in order to fine-tune the output of a pre-trained model, for example, by encouraging certain responses or types of output. The finetuning can be done by either learning a user reward model over user preference data, or by using the preference data directly (through direct preference optimization (Rafailov et al., 2024)). In either case, accurately approximating user preferences is an important task, which becomes way more challenging when the target group of users is heterogeneous (Figure 1) (Pollak and Wales, 1992; Boxall and Adamowicz, 2002).
This paper contributes to this literature by providing a holistic study of learning (different) reward models from heterogeneous user preference data. There are two major challenges in this context. The first (C1) is a pure learning one: preference data from each individual might not be sufficiently rich to construct an accurate model of heterogeneous users. The second (C2) is after learning different reward models for heterogeneous users, how to aggregate them carefully to learn a single model. Moreover, with humans (who are oftentimes viewed as rational decision-makers) involved in the loop, they might strategically misreport their preferences to manipulate this aggregated model. For example, in online rating systems, users may provide extreme feedback to disproportionately influence the overall ratings toward their viewpoint. Our approach develops ways of tackling these challenges.
@@
To address (C1), we adopt two approaches based on representation learning, which assume that individual reward functions share a structure through a common representation. We model each reward function as the inner product of a common representation and a parameter vector. Given the lack of sufficient individual feedback, having a shared structure by representation helps articulate each user’s reward model. The first approach constructs a personalized reward model for each user. In this approach, we find a common representation and learn each individual’s parameter vector by pooling every individual feedback. The second approach segments user preferences into clusters and learns a reward model for each cluster. This approach is useful when individual reward functions might not be available due to insufficient data. By assuming “diversity of user’s parameter vectors”, which means that individual parameter vectors span the entire space of parameters (a common assumption in multi-task learning), we show that this approach enables better sample complexity results. Leveraging data from all users helps learn the common representation, as the diversity assumption guarantees sufficient information about every dimension of the representation.
To address (C2), we first estimate the parameters for each individual’s reward model using the individual’s preference comparison data. Then, we aggregate reward models using a family of reward aggregation rules, which follows six pivotal axioms from social choice theory. We then provide sample complexities of the policy induced from the single aggregated reward model. We additionally provide a model with a different feedback type - probabilistic opinion. Concretely, instead of choosing a single answer from a pool of candidate answers, we allow the human labeler to choose a probability distribution over the answers, which indicates how much the labeler likes those answers. This type of feedback can arguably express the labeler’s preference more accurately. Moreover, probabilistic opinion feedback does not require the relationship between the human reward model and preference. We consider various aggregation rules to aggregate their probabilistic opinion vectors into one. We showed that our suggested probabilistic opinion aggregation rule is equivalent to reward aggregation rules following six pivotal axioms, under the Plackett-Luce model (Plackett, 1975; Luce, 2005).
To deal with the strategic misreport problem, we adopt a mechanism design approach whereby users correctly reporting their preferences is incentivized. We model each human labeler’s utility as a quasi-linear function, considering both the distance between her probabilistic opinion vector and the aggregated opinion vector, and the associated costs. Under this model, we show that our proposed aggregation rule maximizes social welfare. Lastly, we design an incentive-compatible mechanism to guarantee truthful reporting by inducing proper cost in the human feedback collection process.
1.1 Related Works
@@
Empirical evidence has demonstrated the efficacy of incorporating human preferences into reinforcement learning (RL) for enhancing robotics (Abramson et al., 2022; Hwang et al., 2024) and for refining large-scale language models (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). These human inputs take various forms, such as rankings (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022), demonstrations (Finn et al., 2016), and scalar ratings (Warnell et al., 2018). A few approaches have been explored empirically to personalize RLHF. For example, assigning fine-grained rewards to small text segments to enhance the training process (Wu et al., 2024), or training each human labeler’s reward model with Multi-Objective Reinforcement Learning perspective (Jang et al., 2023; Hwang et al., 2024) have been proposed. Moreover, (Li et al., 2024) suggested the training of each human labeler’s reward model directly using personalized feedback with human embedding obtained by the human model, and also an approach for the clustering with finding cluster embedding.
On the theory front, the studies of RLHF have received increasing research interest. The most related prior works are (Zhu et al., 2023; Zhan et al., 2023; Wang et al., 2024), where (Zhu et al., 2023) investigated the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952) within the context of a linear reward framework; while (Zhan et al., 2023) generalized the results to encompass more general classes of reward functions. Both works concern the setting with offline preference data. Additionally, (Kim et al., 2024) provided a linear programming framework for offline reward learning. (Xiong et al., 2024) provided a theoretical analysis for KL-regularized RLHF. In the online setting, (Wang et al., 2024) established a correlation between online preference learning and online RL through a preference-to-reward interface.
Yet, to the best of our knowledge, there is no prior work that has analyzed RLHF with heterogeneous feedback with theoretical guarantees (except the recent independent works discussed in detail below).
@@
Early work of (Baxter, 2000) established a generalization bound that hinges on the concept of a task generative model within the representation learning framework. More recently, (Tripuraneni et al., 2021; Du et al., 2021) demonstrated that, in the setup with linear representations and squared loss functions, task diversity can significantly enhance the efficiency of learning representations. Moreover, (Tripuraneni et al., 2020) provided a representation learning with general representation and general loss functions. Representation learning has been extended to the reinforcement learning setting as well. For low-rank Markov Decision Processes, where both the reward function and the probability kernel are represented through the inner products of state and action representations with certain parameters, (Agarwal et al., 2020; Ren et al., 2022; Uehara et al., 2021) explored the theoretical foundations for learning these representations. Also, (Ishfaq et al., 2024; Bose et al., 2024) analyzed the sample complexity of multi-task offline RL.
@@
Preference aggregation is the process by which multiple humans’ preference orderings of various social alternatives are combined into a single, collective preference or choice (List, 2013). Arrow’s Impossibility Theorem demonstrates that no aggregation rule for preference orderings can simultaneously meet specific criteria essential for ensuring a fair and rational aggregation of each human user’s preferences into a collective decision (Arrow, 1951). Therefore, people considered replacing preference orderings with assigning real numbers to social alternatives (Sen, 2018; Moulin, 2004), which is sometimes called a reward (welfare) function in social choice theory. for each human user. (Skiadas, 2016; Moulin, 2004) provided reward (welfare) aggregation rules which satisfy several desirable properties. Furthermore, an alternative method to circumvent Arrow’s impossibility theorem involved aggregating preferences via probabilistic opinion (Stone, 1961; Lehrer and Wagner, 2012). In this approach, opinions are represented as probability assignments to specific events or propositions of interest.
@@
More recently, (Zhong et al., 2024), which is a concurrent work with this paper, provided a theoretical analysis of reward aggregation in RLHF, focusing primarily on linear representations. Our work, in comparison, considers general representation functions and general relationships between reward function and preference. Unlike (Zhong et al., 2024), where they focused on reward aggregation, we focus on personalization for every human labeler and also employ clustering techniques for personalization. (Zhong et al., 2024) and our paper also both investigated the case that reward and preference are not related. Our paper suggested a probabilistic opinion pooling with a mechanism design to effectively elicit truthful human preferences, presuming human labelers may be strategic. In contrast, (Zhong et al., 2024) analyzed an algorithm for a von Neumann winner policy, where a von Neumann winner policy is a policy that has at least a 50% chance of being preferred compared to any other policy. Moreover, (Zhong et al., 2024) also explored the Pareto efficiency of the resulting policy.
@@
Notation.
The matrix $\mathbf{O}$ denotes an all-zero matrix, while $I$ stands for an identity matrix, of proper dimensions. We use $A\succ\mathbf{O}$ to denote that matrix $A$ is a positive definite matrix. The function $\sigma$ represents the Sigmoid function, defined by $\sigma(x)=1/(1+\exp(-x))$. The notation $[K]$ denotes the set $\{1,2,\dots,K\}$. $\Delta(\mathcal{A})$ refers to a probability vector in $\mathbb{R}^{|\mathcal{A}|}$. The term $\sigma_{k}^{2}(A)$ denotes the $k$-th largest singular value of matrix $A$. A function $f(x)$ is categorized based on the complexity notation as follows: $f(x)=O(g(x))$ if there exists $C>0$ and $x_{0}$ such that $f(x)\leq Cg(x)$ holds for all $x\geq x_{0}$, $f(x)=\Omega(g(x))$ if there exists $C>0$ and $x_{0}$ such that $f(x)\geq C\cdot g(x)$ for all $x\geq x_{0}$, $f(x)=o(g(x))$ if $\lim_{x\to\infty}\frac{f(x)}{g(x)}=0$, and $f(x)=\widetilde{O}(g(x))$ if $f(n)=O(g(x)\cdot\log^{k}(g(x)))$ for some finite $k$. The vector $e_{1}$ is defined as the standard basis vector of proper dimension with the first component being $1$. For a finite-dimensional vector $x$, the norm $\|x\|_{1}$ refers to its $\ell_{1}$-norm, while $\|x\|_{2}$ refers to the $\ell_{2}$-norm, unless otherwise specified. We also define $\|x\|_{\Sigma}=\sqrt{x^{\intercal}\Sigma x}$ for a positive definite matrix $\Sigma$. For a matrix $M$, the norm $\|M\|_{F}$ denotes the Frobenius norm of $M$. The multinomial distribution is denoted by $\text{Multinomial}(p_{1},\ldots,p_{n})$, where $p_{1},\ldots,p_{n}$ are the probabilities of outcomes for each of the $n$ categories, respectively, with $\sum_{i=1}^{n}p_{i}=1$ and $p_{i}\geq 0$ for all $i\in[n]$. Kullback-Leibler (KL) divergence between two probability distributions $P,Q\in\Delta(X)$ is defined as $\sum_{x\in\text{supp}(X)}P(x)\log\left(\frac{P(x)}{Q(x)}\right)$.
2 Preliminaries
Most existing RLHF processes (for language model fine-tuning) consist of two main stages: (1) learning a model of human rewards (oftentimes from preference data), and (2) fine-tuning with the reference policy through Reinforcement Learning algorithms, e.g., Proximal Policy Optimization (PPO) (Schulman et al., 2017). It may also be possible to avoid the explicit learning of reward functions while fine-tuning the policy directly from preference data (Rafailov et al., 2024).
Markov Decision Processes.
We define the state $s$ as an element of the set of possible prompts or questions, denoted by $\mathcal{S}$, and the set of actions $a$, contained in $\mathcal{A}$, as the potential answers or responses to these questions. Consider an RLHF setting with $N$ human labelers (or users), each of whom has their own reward function. This setting can be characterized by a Markov Decision Process (MDP) with $N$ reward functions, represented by the tuple $M=(\mathcal{S},\mathcal{A},H,(P_{h})_{h\in[H]},\bm{r}=(r_{i})_{i\in[N]})$, where $H$ denotes the length of the horizon, $P_{h}:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is the state transition probability at step $h\in[H]$, ${\mathcal{T}}:=(\mathcal{S}\times\mathcal{A})^{H}$ denotes the set of all possible trajectories, and $r_{i}:{\mathcal{T}}\rightarrow\mathbb{R}$ is the reward function for individual $i$ and trajectory $\tau\in{\mathcal{T}}$, representing the utility of human user $i$ from a sequence of responses to a given prompt. We assume $-R_{\max}\leq r_{i}(\tau)\leq R_{\max}$ for every $\tau\in{\mathcal{T}}$ and $i\in[N]$, for some $R_{\max}>0$. This reward model also covers the case that $r_{i}(\tau)=\sum_{h\in[H]}r_{h,i}(s_{h},a_{h})$, where $r_{h,i}:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ denotes the state-action reward function for each step $h$ and individual $i$, and $\tau=(s_{1},a_{1},s_{2},a_{2},\dots,s_{H},a_{H})$. The MDP concludes at an absorbing termination state with zero reward after $H$ steps. A policy $\pi_{h}:(\mathcal{S}\times\mathcal{A})^{h-1}\times{\mathcal{S}}\to\Delta(\mathcal{A})$ is defined as a function mapping trajectories to distributions over actions for each step $h\in[H]$ within the horizon $H$. We define the history-dependent policy class as $\Pi$. The collection of these policies across all steps is denoted by $\pi$$=(\pi_{h})_{h=1}^{H-1}$. The expected cumulative reward of a policy $\pi$ is given by $J(\pi;r_{i}):=\mathbb{E}_{\tau,\pi}[r_{i}(\tau)]$ where the expectation in the formula is taken over the distribution of the trajectories under the policy $\pi$. Trajectory occupancy measures, denoted by $d_{\pi}:{\mathcal{T}}\to[0,1]$, are defined as $d_{\pi}(\tau):=\mathbb{P}_{\pi}(\tau)$, which denotes the probability of generating trajectory $\tau$ following policy $\pi$.
Relationship between Preference and Reward Function.
For the MDP with $M=(\mathcal{S},\mathcal{A},H,(P_{h})_{h\in[H]},\bm{r}=(r_{i})_{i\in[N]})$, if we compare two trajectories $\tau_{0}$ and $\tau_{1}$, we define some random variable $o$ such that $o=0$ if $\tau_{0}\succ\tau_{1}$, and $o=1$ if $\tau_{0}\prec\tau_{1}$. Here, $\tau_{0}\succ\tau_{1}$ indicates that $\tau_{0}$ is preferred than $\tau_{1}$. We assume that $P_{r_{i}}(o=0\mid\tau_{0},\tau_{1})=\Phi(r_{i}(\tau_{0})-r_{i}(\tau_{1}))$ for all $i\in[N]$, where $\Phi:\mathbb{R}\to[0,1]$ is a monotonically increasing function, which satisfy $\Phi(x)+\Phi(-x)=1$ and $\log\Phi(x)$ is a strongly convex function. For example, $\Phi(x)=\sigma(x)$ indicates the BTL model (Definition 4.1 below), a frequently used model for the relationship between preference and reward. Also, we define $P_{\bm{r}}(\cdot\mid\tau_{0},\tau_{1}):=(P_{r_{1}}(\cdot\mid\tau_{0},\tau_{1})^{\intercal},\dots,P_{r_{N}}(\cdot\mid\tau_{0},\tau_{1})^{\intercal})^{\intercal}$. We call $P_{\bm{r}}$ and $P_{r_{i}}$ a preference probability vector induced by the reward vector $\bm{r}$ and the reward $r_{i}$.
3 Provable Personalized RLHF via Representation Learning
3.1 Learning Personalized Reward Model